if a box has 5.0x10^2 J of work done to it, is moved with 25N. What is the distance it moved?
1
Expert's answer
2020-06-16T09:27:04-0400

The work can be calculated as "A = FS\cos\alpha," where S is the distance, "\alpha" is the angle between the directions of force and displacement. We don't know the angle, but we may assume the angle to be equal to 0, so the force and displacement are parallel. So we can calculate the distance as

"S = \dfrac{A}{F} = \dfrac{5.0\cdot10^2\,\mathrm{J}}{25\,\mathrm{N}} = 20\,\mathrm{m}."